A rocket is fired vertically and ascends with a constant vertical acceleration of 20 m/s 2 for 1.0 min. Its fuel is then all used and it continues as a free-fall particle. (a) What is the maxi- mum altitude reached? (b) What is the total time elapsed from take off until the rocket strikes the Earth? (Ignore the variation of g with altitude).

A rocket  is fired  vertically  and  ascends  with  a constant vertical acceleration   of 20 m/s2   for 1.0 min. Its fuel is then all used and  it continues   as  a free-fall   particle.   (a)  What  is the  maxi- mum  altitude  reached?  (b) What  is the total time elapsed  from take off  until  the rocket  strikes  the Earth?  (Ignore  the variation of g with  altitude). 

Grade:12th pass

2 Answers

Sooryakiran
7 years ago
a) s=1/2at^2 =1/2 x 20 x 60 x 60 =36 kmB) t = time to go up + time to reach zero speed under gravity + time to fall free from that altitude.
Soumith Chowdary
7 years ago
Use S=ut+1/2at^2 here S=1/2x20x60x60 S =36000 t=60+60 t=120s because time of ascent =time of descent. u=0